OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 2x + y = 5 x + 3y = 0 A –1 B 1 C 3 D –3

OpenStudy (anonymous):

You need to use elimination to get rid of the x variable and leave only the y. To do that you need to get the coefficient of x as the same value. So multiply the second equation by 2 to get 2x + 6y = 0. Then take that equation away from the first one. The 2x and 2x cancel and y - 6y = -5y. That leaves -5y=5 so y = -1.

OpenStudy (anonymous):

\[2x + y = 5\]\[2x + 6y = 0\]

OpenStudy (anonymous):

you could also use substitution 2x + y = 5 is the same as y = 5 - 2x so then x + 3y = 0 becomes x + 3(5-2x) = 0 or x + 15 - 6x = 0 5x = 15 x=3

OpenStudy (anonymous):

Yeah but make the formula equal to x so you can substitute that in and find y.

OpenStudy (anonymous):

thank you all for your help but i have to give the medal to @alrightatmaths because he had the right answer but thanks again!